OpenStudy (anonymous):

Help? I don't understand this at all... Solve on the interval 0 ≤ θ < 2π Cos(2θ-π/2)=-1

OpenStudy (anonymous):

Basically you are finding values of theta where theta is in the interval 0 <= theta <= 2pi and satisfies the given equation. So basically what the equation says is cos(2theta - pi/2) = -1. So you find all values where cosine is -1 between 0 and 2pi (inclusive). Then these will be the values that 2theta - pi/2 must be equal to because these will be the values where cosine is -1 in the given interval. So you equation 2theta - pi/2 with whatever values you find where cosine is negative one in the given interval, and solve for theta.

OpenStudy (anonymous):

@MoshiBunny If you are completely lost, and don't know thing or don't understand a thing, then I can start you off =)

OpenStudy (campbell_st):

well you need to find the inverse or arc cos then \[2\theta - \frac{\pi}{2} = \cos^{-1}(-1) \] so \[2 \theta - \frac{\pi}{2} = \pi\] you should be able to solve for theta from here..

OpenStudy (anonymous):

\[\cos \left( 2\theta - \frac{ \pi }{ 2 } \right)=-1 |0 \le \theta \le 2\pi \implies 2\theta - \frac{ \pi }{ 2 }=\pi\]Now solve for theta. Can you do that? @MoshiBunny

OpenStudy (anonymous):

So I add pi/2 to both sides?

OpenStudy (anonymous):

& divide by 2?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

What do you get?

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

so 2 theta = 3pi/2?

OpenStudy (anonymous):

ya, now divide by 2.

OpenStudy (anonymous):

theta = 3pi? I feel like that's wrong..

OpenStudy (anonymous):

It is wrong. It's 3pi/4

OpenStudy (anonymous):

\[2\theta = \frac{ 3\pi }{ 2 } \rightarrow \theta = \frac{ 3\pi }{ 2 } \div 2 = \frac{ 3\pi }{ 2 } \times \frac{ 1 }{ 2 }=\frac{ 3\pi }{ 4 }\]You probably made a division mistake. @MoshiBunny

OpenStudy (anonymous):

Yeah, I forgot I had to do the reciprocal ..

OpenStudy (anonymous):

Ya. So you get how the problem was solved? You are basically looking for where the cosine graph is -1 in the interval [0, 2pi]. Graphing it will show that only at pi cosine is -1. That means that the 2theta - pi/2 inside the brackets must equal to pi. So you equate it as 2theta - pi/2 = pi and solve for theta. You get what was done here to get the answer? @MoshiBunny

OpenStudy (anonymous):

So, it would be theta=3pi/4+2kpi?

OpenStudy (anonymous):

No not k. Because we are given an interval to solve for theta for and that interval is [0,2pi], then we only suggest those values of theta in that interval which satisfy the equation. If we weren't told to find theta in a certain interval, only then would we include 'k'. @MoshiBunny

OpenStudy (anonymous):

So it would be just theta = 3pi/4

OpenStudy (anonymous):

Oh.. so there aren't any, like "extra answers"?

OpenStudy (anonymous):

Exactly. The answer with 'k' in it is a general solution, meaning it gives solutions over a domain. Since we are given an interval, we only solve for values satisfying the equation in that interval. So IN THIS CASE, there is no extra answers, just one.

OpenStudy (anonymous):

Thank you! I finally get this :D

OpenStudy (anonymous):

Fan me? =D

OpenStudy (anonymous):

Sure :]

OpenStudy (anonymous):

Btw, if the question has it equal to 0 instead of -1, in the interval [0, 2pi], how many values of theta would we get?

OpenStudy (anonymous):

@MoshiBunny

OpenStudy (anonymous):

1? .. uh.. hehe....

OpenStudy (anonymous):

Nope. We would get 2 answers for theta. This is because in the interval [0, 2pi], the graph of cosine is 0 at two different points, at pi/2 and 3pi/2. So then we solve the inner bracket of 2theta - pi/2 = pi/2 and 2theta - pi/2 = 3pi/2 and that will give us 2 different answers for theta which satisfy the equation in the interval [0,2pi]. Get it? You should really look at the graph of cosine and sine to get a good understanding of this. @MoshiBunny

OpenStudy (anonymous):

Oh.. I looked at my unit circle and I didn't see 3pi/2 on it so I said 1 answer. :O